OpenStudy (anonymous):

sec^2x+2sec=0 solve the equation for 0

OpenStudy (callisto):

sec^2x+2secx=0 secx ( secx +2) =0 secx =0 (rejected) or secx+2 =0 secx =-2 1/cosx =-2 cos x = -1/2 x = 2pi / 3 or x = 4pi/3

OpenStudy (anonymous):

thank you! :)

OpenStudy (callisto):

welcome :)